OpenStudy (anonymous):

Evaluate \[\int\limits_{}^{}\int\limits_{}^{}\int\limits_{E}^{}2xdV\] where E is the region under the plane 2x+3y+z=6 that lies in the first octant.

OpenStudy (anonymous):

i know when z=0, i got the x-y plane graph  but how to know the limit of integration for dz?

OpenStudy (anonymous):

\[ \int _0^3\int _0^{\frac{1}{3} (6-3 x)}\int _0^{-2 x-3 y+6}2 xdzdydx=\frac{9}{4} \]

OpenStudy (anonymous):

owh so we just let the whole equation be in terms of z to integrate z?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thx alot =) will try it out

OpenStudy (anonymous):

yw